Solve the following homogeneous differential equation 2x3 y' = y(2x2 − y2


1
Expert's answer
2021-08-24T08:27:12-0400

"2x^3y'=2x^2y-y^3\\
\frac{dy}{dx}=\frac{2x^2y-y^3}{2x^3}\\
\text{put\,} y=vx\\
v+x\frac{dv}{dx}=\frac{2v-v^3}{2}\\
v+x\frac{dv}{dx}=v-\frac{v^3}{2}\\
x\frac{dv}{dx}=-\frac{v^3}{2}\\
2\frac{dv}{v^3}=-\frac{dx}{x}\\
\frac{-1}{v^2}=-lnx+c\\
v^2=\frac{-1}{c-lnx}\\
\frac{y^2}{x^2}=\frac{1}{lnx+C}\space(put\space C=-c)\\
\text{This is the required solution.}
\\"